OpenStudy (rocklee88):

2. A 5-member committee is to be formed at random from 8 women and 16 men. a) Find the probability that the committee will consist of at least 5 men. b) Find the probability that the committee will consist of at most 3 women.

OpenStudy (sepeario):

a) If there are at least 5 men, then there can be 5,6,7 or 8 men. Therefore 4/(16+8)= 4/24=1/6

OpenStudy (rocklee88):